### v3

# MMSci: A Dataset for Graduate-Level Multi-Discipline Multimodal Scientific Understanding

Abstract
Scientific figure interpretation is a crucial capability for AI-driven scientific assistants built on advanced Large Vision Language Models. However, current datasets and benchmarks primarily focus on simple charts or other relatively straightforward figures from limited science domains. To address this gap, we present a comprehensive dataset compiled from peer-reviewed Nature Communications articles covering 72 scientific fields, encompassing complex visualizations such as schematic diagrams, microscopic images, and experimental data which require graduate-level expertise to interpret. We evaluated 19 proprietary and open-source models on two benchmark tasks, figure captioning and multiple-choice, and conducted human expert annotation. Our analysis revealed significant task challenges and performance gaps among models. Beyond serving as a benchmark, this dataset serves as a valuable resource for large-scale training. Fine-tuning Qwen2-VL-7B with our task-specific data achieved better performance than GPT-4o and even human experts in multiple-choice evaluations. Furthermore, continuous pre-training on our interleaved article and figure data substantially enhanced the model’s downstream task performance in materials science. We will release our dataset to support further research.
Machine Learning, ICML
1 Introduction
Figure 1: The top 20 out of 72 science subjects with the most articles in our dataset MMSci. The corresponding numbers of papers and figures (in brackets) are shown.
Figure 2: Examples of the heterogeneous types of scientific figures in MMSci, collected from open-access, peer-reviewed articles in Nature Communications.
Recent advancements in Large Vision Language Models (LVLMs) (Li et al., 2023; Zhu et al., 2023; Liu et al., 2024; Chen et al., 2024c; Bai et al., 2023b; Achiam et al., 2023; Team et al., 2023; Anthropic, 2024a; Wang et al., 2024a), have demonstrated remarkable capabilities in solving problems involving visual context. The growing capabilities of LVLMs make them promising as AI-driven scientific assistants capable of solving problems and assisting in research in various science domains. A critical aspect of this assistance is interpreting the figures in research articles, which often contain rich, compressed, and complex information, requiring domain-specific expertise to understand.
Current evaluations of LVLMs on scientific figures focus primarily on bar chart interpretation (Kahou et al., 2017; Masry et al., 2022; Roberts et al., 2024; Wang et al., 2024b), and relatively easy figures within limited science domains (Yue et al., 2023, 2024; Li et al., 2024; Chen et al., 2024a). However, figures in scientific articles are far more varied, including microscopy and spectroscopy images, astronomical images, maps, 3D models, molecular structures, geological models, phylogenetic trees, electropherograms, waveforms, heatmaps, spectrograms, etc. Interpreting these figures often requires expert, typically graduate-level, knowledge in specific domains.
To bridge this gap, we introduce MMSci, a comprehensive multimodal dataset curated from open-access Nature Communications articles [^1] https://www.nature.com/ncomms/ under CC BY 4.0 license [^2] https://www.nature.com/ncomms/open-access. The dataset encompasses 72 scientific disciplines, containing 131k articles and 742k figures across diverse visualization types (Figure 2), with discipline distribution shown in Figure 1 (only shows top-20 due to space constraints). To evaluate LVLMs’ comprehension of these complex scientific figures requiring graduate-level expertise, we developed benchmark tasks for figure captioning and multiple-choice questions across different settings.
Our evaluation revealed significant performance gaps among current LVLMs across tasks. For multiple-choice questions, many open-source models performed no better than random guessing. However, some models, such as Qwen2-VL-7B (Wang et al., 2024a) and MiniCPM-V-2.6 (Yao et al., 2024), demonstrated strong performance comparable to proprietary models like Gemini-1.5-Flash (Reid et al., 2024) and Claude-3-Opus (Anthropic, 2024a). GPT-4o (Achiam et al., 2023) and Claude-3.5-Sonnet (Anthropic, 2024b) emerged as the leading models, significantly outperforming other evaluated models. We also conducted human expert evaluations. The results revealed that the leading models achieved performance comparable to or exceeding domain experts, demonstrating their potential as cross-domain scientific assistants. This performance also highlights our tasks’ difficulty and the importance of domain-specific knowledge. While all models struggled with generating precise figure captions, particularly for nuanced semantics, Claude-3.5-Sonnet and GPT-4o still showed markedly improved performance.
Additionally, our dataset provides a vast collection of high-quality research articles and figures across diverse subjects, which can be leveraged as training resources to enhance LVLMs’ understanding of multimodal scientific content. We experimented with constructing visual supervised fine-tuning data, including the task-specific data converted into instruction-following data. This data significantly improved the Qwen2-VL-7B model (Wang et al., 2024a), achieving the highest overall multiple-choice accuracy on our benchmark, though improving captioning performance remained challenging. Furthermore, we pre-trained LVLMs on interleaved article text and figure images, which led to improved performance in material generation, a downstream task in material sciences.
Overall, our contributions are threefold: (1) Data diversity, scope and quality: Our dataset is uniquely composed of high-quality, peer-reviewed academic articles covering 72 diverse scientific disciplines, featuring a wide range of figure types beyond charts. (2) Challenging benchmark: Our benchmark includes tasks with diverse settings to ensure a comprehensive assessment. The evaluation of models and human experts highlights the challenges of the task. (3) Rich training resources: Our dataset provides a valuable training resource. We created task-specific multimodal fine-tuning data and interleaved article and figure data for continuous LVLM pre-training. Our findings highlight the potential of this dataset to improve models’ comprehension of scientific knowledge.
Table 1: Comparison with prior scientific figure understanding benchmark datasets. *The number of subjects in each work is taken from the original paper that uses different taxonomies, offering a sense of the relative coverage in each work.
Benchmark Dataset Data Source Peer-reviewed # Subjects* Image Type Annotations Training Set
FigureQA (Kahou et al., 2017) Synthetic Data N/A N/A Charts Synthetic
DvQA (Kafle et al., 2018) Synthetic Data N/A N/A Charts Synthetic
SciCap (Yang et al., 2023) CS Arxiv Papers 1 (CS) Charts Authentic
SciFiBench (Roberts et al., 2024) CS Arxiv Papers 1 (CS) Charts Authentic
CharXiv (Wang et al., 2024b) Arxiv Papers 8 Charts Human-picked
ArxivCap/QA (Li et al., 2024) Arxiv Papers 32 Open Category Authentic/Synthetic
MMSci (Ours) Nature Communications 72 Open Category Authentic
2 Related Dataset Work
Scientific Figure Understanding.  Scientific figures in academic articles convey rich, valuable information, and there has been extensive research on evaluating their interpretation. As shown in Table 1, existing datasets primarily focus on relatively simple chart figures, which require general chart interpretation skills rather than deep scientific knowledge. Early efforts targeted data visualization figures through synthetic datasets of plots and charts (Chen et al., 2020; Kahou et al., 2017; Kafle et al., 2018). To capture more diverse and complex chart figures, FigureSeer (Siegel et al., 2016) and SciCap (Yang et al., 2023) extracted figures from computer science (CS) papers on arXiv. SciFiBench (Roberts et al., 2024) expanded on SciCap’s chart figures by introducing figure-to-caption and caption-to-figure matching tasks, while CharXiv (Wang et al., 2024b) hand-picked chart figures from arXiv papers. While these datasets focus exclusively on chart figures, ArxivQA/Cap (Li et al., 2024) extended the scope by collecting papers from 32 subjects on arXiv, including various image types beyond charts. However, the collection remains heavily focused on CS and mathematics, with limited coverage of natural sciences. Moreover, since arXiv papers are not peer-reviewed, their quality cannot be guaranteed. In contrast, our dataset comprises peer-reviewed articles from Nature Communications, spanning 72 disciplines and covering a wide range of natural science subjects. We also provide a rich training set for enhancing scientific figure understanding capabilities.
Table 2: The key statistics of MMSci, including the source data and the constructed benchmark test/validation (dev) set and the data for visual fine-tuning in the training set.
Source dataset Number Benchmark test/dev set Number Training set Number
Total subjects 72 Used articles 1,418/1,414 Used articles 128,561
Total articles 131,393 Figure Captioning 1,218 /1,412 Figure Captioning 725,646
Total figures 742,273 Fig2Cap Matching 1,188/1,297 Fig2Cap Matching 84,328
Avg. caption length 153 SubFig2Cap Matching 1,119/1,214 SubFig2Cap Matching 53,882
Avg. figures per article 5.65 SubCap2Fig Matching 1,114/1,221 SubCap2Fig Matching 107,098
Avg. abstract length 150 Multi-turn conversation 108,843
Avg. article length 7,457 Total samples 1,079,797
Multimodal Science Problems.  With the advances in LVLMs, recent studies have focused on evaluating their ability to solve scientific problems involving visual context. However, existing datasets primarily assess models’ ability to ”read” and ”see” simple image content rather than testing their ”understanding” of complex scientific figures. The images in these datasets are relatively straightforward and typically do not require expert scientific knowledge for interpretation. For example, ScienceQA (Lu et al., 2022) focuses on K-12 level problems, while SciBench (Wang et al., 2023) is limited to three disciplines: physics, chemistry, and mathematics. MMMU (Yue et al., 2023) and MMMU-Pro (Yue et al., 2024) cover subjects such as art, business, history, health, humanities, and technology, but their coverage of natural science subjects is limited, and image understanding is not the primary challenge. While MMStar (Chen et al., 2024a) includes natural sciences, its coverage is somewhat limited. In contrast, our work focuses on understanding complex scientific figures that require graduate-level, domain-specific knowledge across scientific disciplines. Our dataset can potentially be used for constructing multimodal science problems, which we leave for future exploration.
3 Data Curation
Source Data Collection.  Our dataset was collected from the Nature Communications website, comprising open-access, peer-reviewed papers across five major categories and 72 subjects. The top 20 subjects are shown in Figure 1, with the full list of all 72 subjects provided in the Appendix, Table 6. Various information regarding each article is easily accessible on this website, providing a user-friendly platform for obtaining all necessary data. For each article, we collected information including the title, abstract, main body content, and references, directly from their respective sections on the article’s webpage (e.g., https://www.nature.com/articles/xxx, where “xxx” is the article’s unique ID). Figures and their captions were obtained from a dedicated figures page under the article’s homepage (e.g., https://www.nature.com/articles/xxx/figures), eliminating the need to extract figures from PDF files and thus ensuring image quality. We used pylatexenc to convert LaTeX expressions of mathematical formulas in the article text and figure captions into plain text. [^3] https://github.com/phfaist/pylatexenc Since these papers are all peer-reviewed and the text, figures, and captions are readily available from the website, ensuring the data is both authentic and high-quality. We thus did not perform additional filtering or content extraction. We crawled articles up to the date of 2024/04/15. The resulting source dataset comprises 131,393 articles and 742,273 figures. More statistics are shown in Table 2.
Sub-caption Extraction.  Many figures in the dataset consist of multiple sub-figures in a single image, with captions that include a main caption and descriptions of each sub-figure (sub-caption), as illustrated in Figure 3. We developed a regular expression matching function to identify sub-figure indices at the beginning of sentences in alphabetical order (a to z), extracting and identifying 514,054 sub-captions/figures, which aids in the consecutive construction of our benchmark.
Figure 3: Illustration of the benchmark data in MMSci. This example is taken from (Guo et al., 2024). The figure (left) contains multiple sub-figures with a main caption (bold) and color-coded sub-captions corresponding to each sub-figure. These sub-figures and sub-captions are used to construct tasks for figure captioning (upper right), sub-figure to sub-caption matching (center right), and sub-caption to sub-figure matching (lower right).
Heterogeneous Figure Types in MMSci We categorized the types of (sub-)figures in MMSci into seven major categories based on a subset of the figures, focusing on the smallest individual components, such as sub-figures when present. Following this manual review, we used GPT-4o to classify the images within the benchmark test set (see benchmark data splits in the next section). Examples of image types are shown in Figure 2, with detailed statistics provided in the Appendix (Section A.1.3). In addition to charts in previous benchmarks, which make up half of the figures, we identified six other major types that vary significantly across different subjects.
4 Benchmarks
We developed two benchmark tasks with varying settings to comprehensively test models’ understanding of scientific figures and content, as shown in Figure 3.
MMSciCap: Scientific Figure Captioning.  Scientific figure captioning in MMSci poses unique challenges compared to natural image captioning. Interpreting figures from Nature Communications articles requires graduate-level domain expertise and understanding of the article’s context. Moreover, these captions are substantially more detailed, averaging 153 words—significantly longer than those in natural image datasets and ArxivCap (Li et al., 2024). This complexity makes our benchmark particularly challenging. We evaluate scientific figure captioning under three settings: (1) Figure-only captioning: Models generate captions solely from the figure without additional context. (2) Abstract-grounded captioning: Models receive both the figure and the paper’s abstract as context. We also evaluated models’ performance when provided with full article content, limiting this assessment to long context proprietary models due to length constraints, detailed in Appendix A.2.2.
For evaluation metrics, we consider overlap-based metrics BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005), the similarity-based metric BERTScore (Zhang et al., 2019), which compare the generated captions to the reference captions, and also the captioning-specific metric CIDEr (Vedantam et al., 2015). Additionally, we use two LLM-based metrics tailored to detailed and complex scientific figure captions: a modified version of FActScore (Min et al., 2023), and G-Eval (Liu et al., 2023b). The modified FActScore breaks down the generated caption $y$ into a set of atomic units, denoted as $\mathcal{A}{y}$. Each atomic unit represents an independent description of either the overall figure or individual sub-figures. We then evaluate whether each atomic unit is supported by the ground-truth caption $\mathcal{C}$. For fine-grained evaluation, the LLM assigns a score $\phi(a,\mathcal{C})$ to each atomic unit $a\in\mathcal{A}{y}$ on a scale from 0 to 1, representing the degree of support from the ground-truth caption. A brevity penalty is incorporated to account for overly concise captions. The overall formulation is defined as follows:
$\begin{gathered}f(y)=\frac{1}{|\mathcal{A}y|}\sum{a\in\mathcal{A}_{y}}\phi(a,\mathcal{C})\cdot\text{exp}(\text{min}(1-\frac{\gamma}{\mathcal{A}_{y}},0)).\end{gathered}$
We set $\gamma$ to 10 in our evaluation. This metric focuses on precision rather than recall. The second metric, G-Eval, compares the generated caption with the reference caption on a scale of 1 to 5, focusing on overall quality.
MMSciQA: Figure Caption Matching.  We construct multiple-choice questions to evaluate models’ ability to match figures with their correct captions across three settings: (1) Figure-to-Caption (Fig2Cap): Models should select the correct main caption from four options, where distractors are drawn from other figures within the same article. This setting tests holistic figure comprehension. (2) Subfigure-to-Subcaption (SubFig2Cap): Given a random sub-figure, models must identify its corresponding sub-caption among four choices drawn from the same figure. This evaluates the ability to interpret specific components within a complex figure. (3) Subcaption-to-Subfigure (SubCap2Fig): In this reverse setting, given a sub-caption, models must select its matching sub-figure from all sub-figures within the same figure. This tests the model’s ability to associate textual descriptions with specific visual elements.
Data Split.  We allocated 1% of articles from each subject to both test and validation sets, yielding 1,418 test and 1,414 validation articles (5-50 articles per subject). Test samples were derived from unique articles to prevent content overlap. For caption tasks, we required a minimum length of 50 words. Each task setting comprised approximately 1,200 samples, balancing coverage and evaluation costs.
5 Training Resources
Our dataset consists of rich articles and figure data, which we explore as training resources to enhance models’ capabilities in comprehending scientific figures and content.
Task-specific Multimodal Training Data.  We created a multimodal training dataset for visual fine-tuning, comprising both single-turn interactions (multiple-choice questions and abstract-grounded figure captioning) and multi-turn conversations about figure interpretation. The multi-turn conversations were generated by transforming figure captions into question-answer pairs about specific sub-figures, with diverse conversation templates generated by GPT-4. All responses were derived from original article content to ensure data quality. This approach yielded over 1 million training instances across 108,843 conversations. Fine-tuning Qwen2-VL-7B (Wang et al., 2024a) on this dataset for one epoch resulted in Qwen2-VL-7B-MMSci.
Table 3: Performance on scientific figure captioning. B2, RL, M, BS, CD, FS, and GE denote BLEU-2, ROUGE-L, METEOR, BERTScore, CIDEr, FActScore, and G-Eval, respectively. *The LLM-based evaluation results, using GPT-4o, are reported on a randomly selected subset of 200 samples. The best results are bolded, with the second-best underlined.
Model Image-only Captioning Abstract-grounded Captioning
B2 RL M BS CD FS* GE* B2 RL M BS CD FS* GE*
Open-source Models
Kosmos2 4.94 11.69 14.53 77.51 0.97 0.87 1.12 2.90 11.81 19.54 79.09 1.62 3.99 1.39
LLaVA1.5-7B 3.15 12.56 11.80 79.93 0.17 3.89 1.08 3.70 13.97 14.54 81.20 0.76 9.07 2.02
LLaVA1.6-Mistral-7B 2.8 10.97 20.45 79.53 0.08 5.17 1.23 3.90 12.70 21.49 80.84 0.48 7.67 1.47
Qwen-VL-7B-Chat 10.02 14.78 15.34 81.95 1.43 3.06 1.28 8.80 15.55 16.02 81.87 2.78 9.14 1.64
InternVL2-2B 1.69 9.60 17.74 78.89 0.03 5.99 1.76 2.27 11.74 18.45 80.88 0.96 10.38 2.17
InternVL2-8B 2.50 11.39 21.07 79.41 0.00 8.01 2.63 3.74 12.30 22.66 80.57 0.02 9.98 3.00
InternVL2-26B 4.18 13.26 24.21 81.02 0.19 12.43 3.01 5.21 14.92 23.19 80.27 2.30 12.31 3.20
IDEFICS2-8B 6.18 9.40 6.51 80.30 0.21 2.56 1.40 6.96 10.81 8.06 80.30 0.65 5.17 1.96
IDEFICS3-8B-Llama3 1.85 10.11 19.09 78.65 0.00 7.26 1.71 2.33 11.28 20.61 79.42 0.15 7.71 1.98
MiniCPM-V-2.6 4.75 14.57 24.84 81.19 1.42 11.15 2.96 6.11 15.36 25.09 82.68 3.27 12.93 2.95
Llama3.2-11B-Vision 2.68 12.98 21.21 78.89 0.08 8.27 2.46 2.60 11.24 22.63 79.63 0.00 9.55 2.18
Qwen2-VL-2B 3.45 12.74 21.39 80.03 0.38 9.94 2.31 5.68 14.47 21.77 81.23 1.43 11.88 2.64
Qwen2-VL-7B 3.60 12.96 23.88 80.06 0.00 10.03 3.39 4.73 14.45 26.00 81.21 0.19 10.36 3.45
Qwen2-VL-7B-MMSci 19.13 21.78 20.89 84.33 3.73 18.17 3.21 21.77 23.46 23.23 84.92 5.94 20.59 3.54
Proprietary Models
Gemini-1.5-Flash 4.84 15.49 26.82 81.10 0.08 8.18 3.70 5.24 16.03 28.71 81.80 0.00 10.14 4.08
Gemini-1.5-Pro 5.40 16.38 27.06 81.13 0.19 14.59 3.79 5.30 16.89 28.91 81.93 0.00 13.76 4.08
Claude-3.5-Sonnet 5.01 15.54 26.32 81.76 0.65 9.39 3.53 5.94 16.65 27.52 81.76 0.46 12.11 4.04
GPT-4V 4.97 14.86 26.62 81.75 0.37 14.17 3.69 5.24 15.65 27.62 82.37 0.20 19.52 4.13
GPT-4o 4.93 15.59 27.02 81.11 0.27 13.20 4.01 5.57 16.36 28.37 81.84 0.36 18.87 4.22
Interleaved Text and Image Data for Pre-training.  MMSci includes full article content and figures, naturally forming interleaved text and image data suitable for pre-training LVLMs (Lin et al., 2023). We discuss the utilization of this interleaved data in Section 7.
6 Benchmark Evaluation Results
Evaluated Models.  We evaluated a range of open-source and proprietary LVLMs, including Kosmos-2 (Peng et al., 2023), Qwen-VL-7B-Chat (Bai et al., 2023a), Qwen2-VL-2B, and Qwen2-VL-7B (Wang et al., 2024a), the LLaVA1.5 and LLaVA-NeXT(1.6) models (Liu et al., 2024, 2023a), IDEFICS2 (Laurençon et al., 2024b) and IDEFICS3 (Laurençon et al., 2024a), the InternVL2 series (Chen et al., 2024b), and Llama3.2-11B-Vision (Team, 2024). For proprietary models, we evaluated Gemini-1.5-Flash and Gemini-1.5-Pro (Reid et al., 2024), Claude-3-Opus (Anthropic, 2024a), Claude-3.5-Sonnet (Anthropic, 2024b), GPT-4V, and GPT-4o (Achiam et al., 2023). The exact model versions used are detailed in Appendix A.2.1.
Table 4: Accuracies (%) of models and human experts on multiple-choice questions. Setting I, II, and III denote Fig2Cap, SubFig2Cap, and SubCap2Fig, respectively.
Model I II III Avg.
Open-source Models
Kosmos2 23.99 23.95 24.33 24.09
LLaVA1.5-7B 32.74 24.31 22.80 26.75
LLaVA1.6-Mistral-7B 34.76 20.38 24.15 26.60
Qwen-VL-7B-Chat 39.56 19.93 27.83 29.23
InternVL2-2B 42.76 33.07 38.42 38.18
InternVL2-8B 52.78 49.60 40.13 47.62
InternVL2-26B 50.59 57.82 71.63 59.81
IDEFICES2-8B 48.65 25.83 21.10 32.21
IDEFICES3-8B-Llama3 50.42 28.43 29.98 36.57
MiniCPM-V-2.6 53.20 58.27 61.67 57.61
Llama3.2-11B-Vision 54.97 45.04 71.18 57.00
Qwen2-VL-2B 60.61 37.62 55.12 51.30
Qwen2-VL-7B 66.16 73.10 79.80 72.87
Qwen2-VL-7B-MMSci 81.48 88.47 92.91 87.48
Proprietary Models
Gemini-1.5-Flash 54.77 77.84 64.41 65.24
Gemini-1.5-Pro 62.79 81.41 77.16 73.52
Claude-3-Opus 52.19 53.17 60.23 55.13
Claude-3.5-Sonnet 68.77 85.34 87.16 80.18
GPT-4V 60.43 75.07 76.12 70.45
GPT-4o 67.42 87.40 84.65 79.57
Random Guess 25.86 24.63 20.62 23.24
PhD Experts 64.18 71.64 72.72 69.51
Scientific Figure Captioning Results.  As shown in Table 3, grounding captions in article abstracts consistently improves generation quality across all models by providing essential context. While most models struggle to capture the nuanced semantics and style of ground truth captions, our fine-tuned model effectively learned these subtle details from the training data. Although Qwen-VL-7B-Chat achieves high scores on certain metrics, this is primarily due to its tendency to generate concise outputs. The consistently low CIDEr scores across models highlight the distinct challenges of captioning our figures than natural images.
In terms of LLM-based metrics, which evaluate quality beyond semantic nuance, open-source models significantly underperform compared to proprietary models. For G-Eval, which assesses overall similarity to reference captions, proprietary models achieve superior performance. However, on FActScore, which measures precision in describing specific figure components, our fine-tuned model performs best. Nevertheless, all models fall short of satisfactory performance, highlighting the ongoing challenge of precise scientific figure description.
Multi-choice Question Results.  Table 4 presents the multiple-choice question results across three settings. The Figure-to-Caption (Setting I) task, requiring models to identify correct summaries of multi-panel figures (examples in Figure 7, Appendix), proved most challenging. Our fine-tuned model outperformed the strongest proprietary model by over 10%. For SubFig2Cap (Setting II) and SubCap2Fig (Setting III), proprietary models significantly outperformed open-source models, suggesting limitations in identifying nuanced figure content. While some open-source models (LLaVA1.5, LLaVA1.6, Qwen-VL-7B-Chat) performed at random-chance levels, others (MiniCPM-V-2.6, Llama3.2-11B-Vision, Qwen2-VL-7B) demonstrated strong competitiveness. Although Claude-3.5-Sonnet and GPT-4V led among proprietary models, our fine-tuned Qwen2-VL-7B-MMSci achieved the highest overall performance, validating our training data’s effectiveness.
PhD Expert Evaluations.  We organized the dataset into 10 major scientific categories aligned with the Prolific platform [^4] https://www.prolific.com/: Material Science, Chemistry, Physics, Biochemistry, Environment, Climate Sciences, Earth Sciences, Biological Sciences, Biomedical Sciences, and Health and Medicine. For each category, we selected 75 questions (25 per setting) and recruited three PhD-level evaluators with verified degrees in each domain through Prolific, totaling 30 experts. The evaluators provided two assessments: (1) Question Quality Assessment: Experts rated question clarity and domain-knowledge testing effectiveness on a 5-point scale; and (2) Human Expert Performance: Experts answered questions with a suggested time limit of one minute to establish a human performance baseline.
The 30 PhD experts gave an average question quality score of 4.01 (4.09 for Fig2Cap, 4.03 for SubFig2Cap, and 3.91 for SubCap2Fig), where a score of 4 indicates that the question is clear, answerable, and requires an adequate understanding of the scientific content in the figure. This validates the quality of our benchmark questions. The averaged highest-performing expert results for each domain are reported in Table 4. Notably, our fine-tuned model and leading proprietary models achieved performance surpassing PhD-level experts, who were constrained to one minute per question. This superior performance likely reflects the models’ ability to rapidly process dense scientific information across various domains, highlighting both the task’s complexity and the potential of LVLMs as efficient cross-domain scientific assistants. Detailed evaluation procedures and results are provided in Appendix A.2.3.
7 A Case Study in Material Sciences
Material science as the subject with the most articles and figures in our dataset, is an important and highly interdisciplinary field that requires knowledge from various subjects. Given its significance, we conducted a case study to explore how our dataset could enhance material science knowledge. Previous research has investigated the application of language models to material science tasks (Walker et al., 2021; Rubungo et al., 2023; Miret & Krishnan, 2024). A recent study (Gruver et al., 2024) demonstrated promising results using LLaMA2 (Touvron et al., 2023) for material generation by representing crystal structures as text strings and training the model to generate these structures. However, LLaMA2’s scientific knowledge may be insufficient for fully understanding material generation principles. To address this limitation, we explored continuous pre-training of LLaMA2 using our interleaved scientific article and figure dataset, aiming to improve the model’s performance on stable material generation tasks.
Visual Pre-Training on MMSci We continuously pre-trained the LLaMA2-7B model on our collected interleaved article text and figure images, using data within materials science as well as other eight related subjects in the same Physical Science category. To achieve that, we leverage LLaVA’s architecture (Liu et al., 2024), equipping LLaMA2 with a pre-trained CLIP ViT-L/14-336 (Radford et al., 2021) as the visual encoder and a 2-layer MLP as the projector. During training, we initially kept the LLM frozen and used data from general domains provided by (Liu et al., 2024) to initialize the projector. We then trained the model on the interleaved text and image data from general domains in MMC4 (Zhu et al., 2024) to further develop its image perception abilities, followed by our collected interleaved articles and figures in MMSci to infuse scientific knowledge. In this stage, we tuned both the LLM and the projector, for one epoch. For the resulting multimodal model, we use its LLM part, named LLaMA2-7B-MMSci, for the subsequent material generation.
Figure 4: The prompt for generating crystal structure.
Fine-tuning for Materials Generation.  Given the LLM, we further fine-tune it for the material generation task as in (Gruver et al., 2024). Specifically, periodic materials are characterized by a unit cell that repeats infinitely in all three dimensions. Each unit cell is specified by its side lengths ($l_{1}$, $l_{2}$, $l_{3}$) and angles ($\theta_{1}$, $\theta_{2}$, $\theta_{3}$). Within this lattice structure, there are $N$ atoms, each identified by an element symbol, $e_{i}$, and a set of 3D coordinates ($x_{i}$, $y_{i}$, $z_{i}$). Tzhe structure of a bulk material $C$ can be represented by:
$C=(l_{1},l_{2},l_{3},\theta_{1},\theta_{2},\theta_{3},e_{1},x_{1},y_{1},z_{1},...,e_{N},x_{N},y_{N},z_{N}).$
The prompt for generating these structures is shown in Figure 4. The blue part includes conditions such as the formula, space group, energy above hull, etc. The red part is the generated representation of the crystal structure, and the text above is the prompt.
Following (Xie et al., 2021; Gruver et al., 2024), we use the MP-20 dataset (Jain et al., 2013) of 45,231 stable materials, where successful generation should produce at least metastable crystals. The training data incorporates both conditional generation prompts (single or multiple conditions) and infilling prompts for masked crystal structure strings. Training is limited to one epoch to maintain diversity in generated materials.
Table 5: Evaluation of unconditional material generation covering validity, coverage and property distribution, and stability checks. Performance reported over 10,000 samples.
Method Validity Check Coverage Property Distribution Metastable Stable
[…]
Previous non-language baselines
CDVAE 1.000 0.867 0.992 0.995 0.688 1.432 22.1% 1.2%
LM-CH 0.848 0.836 0.993 0.979 0.864 0.132 N/A N/A
LM-AC 0.958 0.889 0.996 0.986 0.696 0.092 N/A N/A
GPT-4o with Few-shot Prompting
GPT-4o 5-shot 0.799 0.898 0.280 0.961 5.421 1.017 1.50% -
GPT-4o 10-shot 0.787 0.820 0.654 0.963 3.976 0.917 4.72% 0.09%
Gruver et al. (2024): LLaMA2 with Task-specific Fine-Tuning
LLaMA2-7B 0.967 0.933 0.923 0.950 3.609 1.044 33.6% 2.1%
LLaMA2-13B 0.958 0.923 0.884 0.983 2.086 0.092 34.3% 4.9%
LLaMA2-70B 0.997 0.949 0.860 0.988 0.842 0.433 50.1% 5.3%
Ours: LLaMA2 with Continuous Pre-Training on MMSci plus Task-specific Fine-Tuning
LLaMA2-7B-MMSci 0.993 0.979 0.916 0.996 1.675 0.353 64.5% 8.2%
Fraction of structures that are first predicted by M3GNet to have $E_{\text{hull}}^{\text{M3GNet}}<0.1$ eV/atom, and then verified with DFT to have $E_{\text{hull}}^{\text{DFT}}<0.0$ eV/atom.
Results.  For evaluation, we perform unconditional generation of potential stable materials using a temperature of 0.7 to sample 10,000 structures (Xie et al., 2021; Gruver et al., 2024). We assess performance across four metrics: validity (adherence to physical constraints), coverage and property metrics (alignment with ground truth distribution), and stability (percentage of samples deemed metastable by M3GNet (Chen & Ong, 2022) and stable by DFT (Hafner, 2008)). As shown in Table 5, GPT-4o fails at this task without specific training. LLaMA2-7B achieves superior results after continuous pre-training on our interleaved articles and figures, followed by multi-task fine-tuning. This model demonstrates the best performance in compositional validity, coverage precision, and the most important metrics metastability, and stability, highlighting the benefit of our data in enhancing the generative model’s acquisition of scientific knowledge.
Figure 5: Ablation studies on the influence of different pre-training data over LLaMA2-7B.
Ablation Studies.  To understand the factors contributing to LLaMA2-7B-MMSci’s performance, we explored different pre-training data configurations: using only interleaved data from either MMC4 (general interleaved data) or MMSci, using interleaved data from MMC4 combined with text-only data from MMSci, and using no additional pre-training data, followed by the same fine-tuning setup. As shown in Figure 5, the text-only and interleaved data from MMSci achieved the top-2 overall performance when combined with MMC4 which equips the model to effectively read text and interpret images within scientific articles. Using both articles and figures led to better performance than using text-only data from MMSci, highlighting the importance of understanding both figures and content in scientific literature. In contrast, using only general domain data from MMC4 did not result in improvements, and directly training on MMSci even slightly decreased performance in structure validity. This is likely because incorporating visual information can confuse the model if it has not been sufficiently pre-trained with general interleaved data. Overall, continuous pre-training on our data shows the potential to infuse scientific knowledge that enhances downstream tasks.
8 Conclusion
In this work, we present MMSci, a multidisciplinary multimodal dataset containing high-quality, peer-reviewed articles and figures across 72 scientific disciplines. Using this dataset, we construct a challenging benchmark to evaluate the capabilities of LVLMs in understanding scientific figures and content, revealing significant deficiencies. Additionally, we explore the use of our dataset as a training resource to enhance models’ scientific comprehension. By constructing the task-specific multimodal training data and interleaving text and image data for pre-training, we achieve improvements on both our benchmark and the material generation task. Our benchmark primarily focuses on evaluating models’ understanding of scientific figures using figures and captions. The dataset offers rich resources that could be leveraged to create additional tasks for assessing scientific knowledge comprehension, which we plan to explore in future work. Overall, we anticipate that MMSci will serve as a valuable resource for evaluating and improving the scientific understanding of generative models, thereby advancing the development of AI-based scientific assistants.
Impact Statement
Our dataset provides significant benefits by serving as an evaluation benchmark for assessing large multimodal models’ understanding of scientific articles and figures, as well as a training resource to enhance their performance in scientific and research-related tasks. There might be potential societal consequences, like potential misuse in academic integrity, such as academic fraud or improper assistance when using the models trained with our data. However, this is a widespread issue in the current era of large language models and is not limited to our work.
Appendix A Appendix
A.1 Dataset Description
A.1.1 Data and Code Access
We provide access to our data, model checkpoints, and code through the following links:
Figure 6: The five major categories and 72 subjects in our dataset.
A.1.2 Subjects
Our dataset spans five major categories and includes 72 distinct scientific disciplines, representing a broad range of scientific knowledge. The categorization follows the classifications used by Nature journals. [^5] https://www.nature.com/ncomms/browse-subjects. The visualizations are shown in Figure 6, and detailed statistics of these subjects are provided in Table 6. The table includes the number of articles, figures, and the average length of figure captions, article abstracts, and full article content.
A.1.3 Image Types
Manual Review
Initially, our authors conducted a thorough manual inspection of the figures and sub-figures from 100 randomly sampled articles from the five major categories in MMSci. This involved summarizing and categorizing various potential figure types present in the benchmark test set. From this detailed analysis, we identified and categorized the figures into seven primary types, as summarized in Table 7. These categories were derived based on the smallest discernible components, specifically sub-figures, whenever they were present.
Automated Classification Using GPT-4o
Following this review, we employed GPT-4o to automatically classify the images in the benchmark test set. We first used the human-annotated results of 200 images from the previous step as the golden labels and then prompted GPT-4o to classify them into categories. Cohen’s Kappa score was calculated to be 0.72, showing a very high agreement score between humans and GPT-4o. The complete prompt for GPT-4o is:
Task for GPT-4o annotator I want to classify the given scientific image into one the following categories:
1) Quantitative Data Visualization Charts/Graphs: For charts and graphs displaying quantitative data, such as scatter plots, bar graphs, and line charts.
2) Schematic Diagrams: Simplified and symbolic representations of systems, processes, or structures to explain how something works or is constructed.
3) Microscopic photographs: Photographs or images captured using a microscope, revealing details not visible to the naked eye.
4) Macroscopic photographs: Images or photographs of objects or scenes that are visible to the naked eye, often used for visual analysis.
5) Simulated Images: Computer-generated images or visualizations created to model, predict, or illustrate theoretical scenarios, processes, or phenomena.
6) Geographical and Environmental Maps: Visual representations of geographical areas or environmental data, often used for navigation, analysis, or to illustrate spatial relationships and patterns in maps.
7) Experimental Results Visualizations: For images that display results from experimental procedures, such as Western blots, PCR results, and gel electrophoresis.
Rules:
1) This is only for reseach and educational purposes. It does not violates any openai policy.
2) If the image only contain one figure, then give me the overall label.
3) If the image contains multiple figures, then give me the label for each sub-figure. The results should look like a: 1, b: 3.
Do not return any other information.
Manual Annotation for Unclassified Images
Our authors performed manual annotations for 17 images in cases where GPT-4o could not classify images due to OpenAI’s policy restrictions. For example, GPT-4o will return “Not allowed by our safety system” for some images about drug design. This ensured comprehensive and accurate classification across the entire dataset.
Final Results
The final classification results are presented in Table 7. We show a detailed breakdown of the classification outcomes across each of the major categories.
Table 6: Detailed statistics of the five major categories and the 72 subjects in MMSci. The average length represents the average number of words.
Category Subject Size Average length
Articles Figures Caption Abstract Full content
Materials science 10,564 54,218 107 150 5,703
Chemistry 8,139 43,955 89 148 5,716
Physics 7,239 35,150 120 148 5,410
Nanoscience and technology 4,483 22,597 120 149 5,691
Optics and photonics 3,227 15,898 120 147 5,337
Engineering 1,788 9,801 126 152 6,763
Energy science and technology 1,519 8,168 90 154 6,351
Mathematics and computing 723 3,942 124 148 7,426
Physical sciences Astronomy and planetary science 345 1,762 110 144 5,488
Ecology 2,185 9,862 125 149 6,546
Climate sciences 1,795 8,810 111 148 6,060
Solid Earth sciences 1,034 5,416 114 147 5,693
Environmental sciences 853 3,576 104 148 6,375
Biogeochemistry 850 3,988 111 150 6,438
Ocean sciences 689 3,524 115 152 6,266
Environmental social sciences 452 2,069 99 145 6,534
Natural hazards 311 1,686 109 141 6,341
Planetary science 406 1,997 109 145 5,549
Hydrology 260 1,258 110 149 6,101
Limnology 65 280 120 146 6,212
Earth and environmental sciences Space physics 126 717 109 146 5,339
Cell biology 6,490 44,111 204 149 8,968
Biochemistry 6,145 37,608 168 149 8,330
Microbiology 5,225 29,487 167 153 7,966
Neuroscience 5,016 32,162 198 148 9,410
Molecular biology 4,843 31,000 193 149 8,955
Genetics 4,665 25,037 169 150 8,165
Cancer 5,215 32,779 196 151 8,820
Immunology 4,024 26,103 195 152 8,781
Biological techniques 3,540 20,169 176 147 8,297
Computational biology and bioinformatics 2,914 16,084 162 150 8,523
Biotechnology 2,633 14,689 170 147 8,118
Biophysics 2,440 14,315 166 150 7,923
Structural biology 3,432 20,402 155 150 8,024
Ecology 2,223 10,052 126 149 6,561
Developmental biology 2,205 14,947 199 151 9,018
Evolution 1,941 9,493 144 150 7,202
Plant sciences 1,659 9,528 163 151 7,846
Physiology 1,619 10,649 190 150 8,892
Chemical biology 1,812 10,523 150 147 7,885
Systems biology 993 5,594 184 149 8,674
Drug discovery 964 5,877 174 150 8,675
Stem cells 1,191 7,870 205 152 9,277
Zoology 502 2,347 144 150 6,613
Biological sciences Psychology 410 2,066 154 148 8,744
Diseases 3,459 20,256 177 152 8,060
Medical research 1,839 10,171 167 154 7,572
Oncology 1,161 7,140 196 156 8,897
Health care 880 4,357 137 150 6,701
Pathogenesis 505 3,223 190 151 8,157
Biomarkers 558 2,959 168 152 7,905
Cardiology 400 2,580 188 152 8,927
Gastroenterology 406 2,670 188 154 8,792
Endocrinology 393 2,590 192 156 9,104
Anatomy 378 2,431 187 147 8,098
Neurology 355 2,164 179 153 8,741
Molecular medicine 342 2,100 187 150 8,697
Risk factors 246 1,058 135 154 6,870
Rheumatology 153 999 191 151 8,969
Nephrology 137 943 193 153 9,194
Signs and symptoms 50 262 169 148 7,270
Urology 38 232 198 155 8,681
Health sciences Health occupations 2 12 84 162 5,666
Social sciences 393 1,713 114 143 6,848
Scientific community 127 363 123 90 4,576
Energy and society 158 827 95 149 6,991
Agriculture 85 396 107 147 6,581
Developing world 75 330 111 128 5,986
Water resources 61 289 100 150 6,531
Geography 49 228 101 144 6,444
Business and industry 46 233 94 143 6,441
Scientific community and society Forestry 43 185 107 148 6,618
Total 72 131,393 742,273 153 150 7,457
Table 7: The figure types in the benchmark test set of MMSci regarding the five major categories, where C1-C5 represents Physical sciences, Earth and environmental sciences, Biological sciences, Health sciences, and Scientific community and society, respectively.
Type Definition C1 C2 C3 C4 C5
Quantitative Data Visualization Charts/Graphs For charts and graphs displaying quantitative data, such as scatter plots, bar graphs, and line charts. 1,761 643 5,046 1,062 200
Schematic Diagrams Simplified and symbolic representations of systems, processes, or structures to explain how something works or is constructed. 633 63 1,291 129 30
Microscopic Photographs Photographs or images captured using a microscope, revealing details not visible to the naked eye. 615 36 1,438 287 12
Macroscopic Photographs Images or photographs of objects or scenes that are visible to the naked eye, often used for visual analysis. 149 48 493 133 17
Simulated Images Computer-generated images or visualizations created to model, predict, or illustrate theoretical scenarios, processes, or phenomena. 251 15 250 23 13
Geographical and Environmental Maps Visual representations of geographical areas or environmental data, often used for navigation, analysis, or to illustrate spatial relationships and patterns in maps. 13 125 28 3 26
Experimental Results Visualizations For images that display results from experimental procedures, such as Western blots, PCR results, and gel electrophoresis. 47 3 1,120 290 1
Total - 3,469 933 9,666 1,927 299
Table 8: Evaluated LVLMs in our experiments with their versions or Huggingface model paths.
Model Model versioning/path
GPT-4V gpt-4-turbo-2024-04-09
GPT-4o gpt-4o-2024-05-13
Gemini-1.5-Pro gemini-1.5-pro-001
Gemini-1.5-Flash gemini-1.5-flash-001
Claude-3.5-Sonnet claude-3-5-sonnet-20240620
Claude-3-Opus laude-3-opus-20240229
Kosmos2 https://huggingface.co/microsoft/kosmos-2-patch14-224
LLaVA1.5-7B https://huggingface.co/llava-hf/llava-1.5-7b-hf
LLaVA1.6-Mistral-7B https://huggingface.co/llava-hf/llava-v1.6-mistral-7b-hf
Qwen-VL-7B-Chat https://huggingface.co/Qwen/Qwen-VL-Chat
InternVL2-2B https://huggingface.co/OpenGVLab/InternVL2-2B
InternVL2-8B https://huggingface.co/OpenGVLab/InternVL2-8B
InternVL2-26B https://huggingface.co/OpenGVLab/InternVL2-26B
IDEFICS2-8B https://huggingface.co/HuggingFaceM4/idefics2-8b
IDEFICS3-8B-Llama3 https://huggingface.co/HuggingFaceM4/Idefics3-8B-Llama3
MiniCPM-V-2.6 https://huggingface.co/openbmb/MiniCPM-V-2_6
Llama3.2-11B-Vision https://huggingface.co/meta-llama/Llama-3.2-11B-Vision-Instruct
Qwen2-VL-2B https://huggingface.co/Qwen/Qwen2-VL-2B-Instruct
Qwen2-VL-7B https://huggingface.co/Qwen/Qwen2-VL-&B-Instruct
A.2 Experimental Setup
A.2.1 Evaluated Model
The exact model versions used are detailed in Table 8. All inferences for the open-source models were executed on a computing cluster equipped with eight NVIDIA A100 GPUs, each with 40GB of memory.
A.2.2 Captioning Evaluation
FActScore Evaluation
We modified the FActScore, which was originally designed to evaluate the factual accuracy of generations using external knowledge sources like Wikipedia. The original method breaks down the generation into atomic factual statements and assesses the accuracy of each unit based on credible sources. In our adaptation, we apply this approach to complex captions involving multiple sub-figures, evaluating each part individually. Since there is no external knowledge source, we assess each atomic unit based on the ground-truth caption. This process involves two steps.
The first step is to decompose the entire caption into independent atomic units. We provide the model with an example for this step, as shown below:
Prompt for Caption Decomposition Your task is to break down the caption into separate, independent descriptions for the entire figure and each panel, formatted appropriately and separated by ’-’.
The figure consists of four sub-figures labeled a, b, c, and d. All four images appear to be scanning electron microscope (SEM) images showing the microstructure of different materials, likely related to the iron-based cathode catalysts described in the article.
a. This image shows a highly porous structure with interconnected particles forming a network. The particles appear to be in the nanometer to micrometer size range. The scale bar indicates 1 $\mu$m. b. This image displays a closer view of what seems to be a similar material to (a), but at a higher magnification. The individual particles are more clearly visible, showing a rough, granular texture. The scale bar indicates 200 nm. c. This image reveals a different morphology compared to (a) and (b). It shows larger, more consolidated structures with a rougher surface texture. There are still visible pores and gaps between the structures. The scale bar indicates 1 $\mu$m. d. This final image appears to be a cross-sectional view of a porous material, possibly showing the internal structure of the catalyst. It reveals a highly interconnected network of pores and channels throughout the material. The scale bar indicates 200 nm.
These images likely represent different stages or variations of the iron-acetate/phenanthroline/zeolitic-imidazolate-framework-derived electrocatalyst mentioned in the article. The varying structures and porosities shown in these images could be related to the enhanced mass-transport properties and increased volumetric activity described in the text.
Independent Descriptions:
- The figure consists of four sub-figures labeled a, b, c, and d.
- All four images appear to be scanning electron microscope (SEM) images.
- The images show the microstructure of different materials.
- The materials are likely related to the iron-based cathode catalysts described in the article.
- Image a shows a highly porous structure with interconnected particles forming a network.
- The particles in image a are in the nanometer to micrometer size range. - The scale bar in image a indicates 1 $\mu$m.
- Image b displays a closer view of a material similar to that in image a but at higher magnification.
- The individual particles in image b are more clearly visible and show a rough, granular texture.
- The scale bar in image b indicates 200 nm.
- Image c shows larger, more consolidated structures with a rougher surface texture.
- There are visible pores and gaps between the structures in image c.
- The scale bar in image c indicates 1 $\mu$m.
- Image d appears to be a cross-sectional view of a porous material.
- Image d reveals the internal structure of the catalyst with a highly interconnected network of pores and channels.
- The scale bar in image d indicates 200 nm.
- These images likely represent different stages or variations of the iron-acetate/phenanthroline/zeolitic-imidazolate-framework-derived electrocatalyst mentioned in the article.
- The varying structures and porosities shown in these images could be related to the enhanced mass-transport properties described in the text.
- The varying structures and porosities in the images may contribute to increased volumetric activity described in the article.
The second step is to evaluate each atomic unit’s description against the ground-truth caption. In this step, we use zero-shot prompting. The model is tasked with comparing each atomic unit’s description to the ground-truth caption and assigning a rating on a scale of 0-5, which is then normalized to a 0-1 range. The prompt is as follows:
Prompt for Atom Unit Description Rating How relevant is the generated caption to the provided human-written caption for the figure? Determine the extent to which the information in the generated caption is included or referenced in the human-written caption. Respond with a score between 0 and 5.
Human-written caption: {REFERENCE}
Generated caption: {GENERATION}
G-Eval Evaluation
Our G-Eval evaluation follows the implementation in (Liu et al., 2023b). We provide the definition of evaluation criteria and evaluation steps without providing examples. The model is tasked with assigning a score in the range of 1-5. The detailed prompt is as follows:
Prompt for G-Eval Evaluation You will be given a oracle caption that describes a figure. You will then be given a second caption written for the same figure. Your task is to rate the second caption on one the following metric.
Evaluation Criteria: Relevance (1-5) - The extent to which the second caption is relevant to the key elements and context described in the oracle caption. A relevant caption should focus on the same subjects, objects, actions, or context highlighted in the oracle caption, without introducing unrelated or extraneous details.
Evaluation Steps: 1. Review the Oracle Caption: Carefully read the oracle caption to understand the main elements and context it describes.
2. Review the Second Caption: Assess whether the second caption focuses on the same key elements and context as the oracle caption. Evaluate if the second caption stays on topic and does not introduce irrelevant details.
3. Assign a Score for Relevance: Based on the Evaluation Criteria, rate how relevant the second caption is to the oracle caption’s description of the same image.
Table 9: Hyperparameters for visual supervised fine-tuning.
Hyperparameter Values
base model https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct
epochs $1$
global batch size 8
learning rate $0.0001$
learning rate scheduler cosine
weight decay $0.0$
warmup ratio $0.1$
max length $4096$
lora modules q_proj, k_proj, v_proj, o_proj, up_proj, gate_proj, down_proj
Table 10: Performance comparison on scientific figure captioning task when grounded on abstract and full article.
Model Context BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L Meteor BertScore FActScore G-Eval
Gemini-1.5-Flash Abstract 3.29 26.74 7.47 16.03 28.71 81.80 10.14 4.08
Gemini-1.5-Pro Abstract 3.33 28.71 7.73 16.89 28.91 81.93 13.76 4.08
Claude-3.5-Sonnet Abstract 3.20 29.60 6.71 16.65 27.52 81.76 12.11 4.04
GPT-4V Abstract 3.18 28.45 7.01 15.65 27.62 82.37 19.52 4.13
GPT-4o Abstract 3.58 28.85 7.79 16.36 28.37 81.84 18.87 4.22
Gemini-1.5-Flash Full Article 6.94 32.83 14.15 22.02 34.50 83.26 19.41 4.12
Gemini-1.5-Pro Full Article 7.01 32.24 13.34 19.32 33.75 83.18 19.33 4.22
Claude-3.5-Sonnet Full Article 7.99 37.63 13.61 23.63 34.66 84.34 21.67 4.52
GPT-4V Full Article 5.65 33.09 10.95 19.25 31.46 83.48 23.18 4.24
GPT-4o Full Article 9.90 37.06 17.63 24.89 37.52 83.64 24.12 4.58
Captioning Grounded on Full Article
We also explored using entire articles as context for captioning. Due to the average article length exceeding 10k tokens, we evaluated this approach only on proprietary models capable of handling long contexts: GPT-4o, GPT-4V, Claude-3.5-Sonnet, and Gemini-1.5-Pro/Flash. As shown in Table 10, providing the full article as context improved performance compared to using only the abstract. This improvement is reasonable since understanding scientific figures typically requires grounding in the article’s content, as abstracts alone may not provide sufficient context. However, we note that this approach may potentially benefit from content repetition, as similar descriptions might appear in both the caption and the article text.
Table 11: Recategorization of the 72 subjects in MMSci dataset for recruiting Phd experts of each major category from Prolific platform.
Re-categoried Fields Original Subjects from Nature Communications
Material Science Materials science, Nanoscience and technology
Chemistry Chemistry
Physics Physics, Optics and photonics
Engineering Engineering
Energy Energy science and technology, Energy and society
Mathematics and Computing Mathematics and computing
Astronomy and Planetary Science Astronomy and planetary science, Planetary science, Space physics
Environment Ecology, Environmental sciences, Biogeochemistry, Water resources
Climate Sciences Climate sciences
Earth Solid Earth sciences, Ocean sciences, Natural hazards, Hydrology, Limnology, Geography
Social Sciences Environmental social sciences, Psychology, Social sciences, Scientific community, Developing world
Biochemistry Biochemistry, Molecular biology, Biophysics, Structural biology, Chemical biology
Biological Sciences Microbiology, Genetics, Biological techniques, Computational biology and bioinformatics, Developmental biology, Evolution, Plant sciences, Physiology, Systems biology, Zoology, Cell biology
Biomedical Sciences Neuroscience, Immunology, Biotechnology, Stem cells, Pathogenesis, Biomarkers, Anatomy, Molecular medicine
Health and Medicine Cancer, Diseases, Medical research, Health care, Oncology, Cardiology, Gastroenterology, Endocrinology, Neurology, Risk factors, Rheumatology, Nephrology, Signs and symptoms, Urology, Health occupations
Pharmacology Drug discovery
Agriculture Agriculture, Forestry
Business and Industry Business and industry
A.2.3 Human Expert Evaluation
To analyze our dataset, we recruited domain experts (PhDs in corresponding fields) through the online professional annotation platform Prolific [^6] https://www.prolific.com/. We refined and consolidated the original 72 subject categories from Prolific into 18 broader groups to balance between comprehensive coverage and sufficient specificity. The recategorized subjects are shown in Table 11. From these 18 recategorized fields, we focused on 10 major scientific domains where PhD annotators were available on Prolific. We recruited 30 PhDs as human evaluators with verified degrees in these domains: Material Science, Chemistry, Physics, Biochemistry, Environment, Climate Sciences, Earth Sciences, Biological Sciences, Biomedical Sciences, and Health and Medicine. Each evaluator provided two types of assessments: Question Quality Assessment and Expert Performance Score. The results for each group are detailed in Table 12.
Question Quality Assessment
For the quality assessment, evaluators were asked to assess whether the questions were clear and demonstrated understanding of scientific knowledge within their respective disciplines. They used the following 5-point scale:
  • Score Point 1: The question is irrelevant or cannot be answered based on the scientific content presented in the figure.
  • Score Point 2: The question lacks clarity or can be answered without specific knowledge of the scientific content in the figure (e.g., it can be answered with common sense).
  • Score Point 3: The question is clear but requires only minimal understanding of the scientific content in the figure.
  • Score Point 4: The question is clear, answerable, and requires an adequate understanding of the scientific content in the figure.
  • Score Point 5: The question is clear, answerable, and effectively evaluates a very deep understanding of the scientific content in the figure.
Expert Performance Score
For the expert evaluation tasks, we created a subset of questions for each category by selecting 25 questions per setting (75 total) from our three figure-caption matching tasks in the original test set. We report the results from the best-performing expert (who achieved the highest average performance) in each category. The annotators were instructed to select their answers within a one-minute time limit per question.
Table 12: Quality scores and Phd experts’ accuracies across the ten re-grouped fields.
Field Fig2Cap SubFig2Cap SubCap2Fig
Quality (1-5) Accuracies (%) Quality (1-5) Accuracies (%) Quality (1-5) Accuracies (%)
Material Science 4.0267 92.00 4.2933 92.00 4.1333 84.00
Chemistry 4.1333 84.00 3.7467 92.00 3.6133 100.00
Physics 4.0267 48.00 3.5467 72.00 3.8133 80.00
Biochemistry 3.1600 80.00 4.8267 56.00 4.4133 72.00
Environment 4.1067 44.00 4.4667 64.00 4.3467 76.00
Climate Sciences 4.1296 77.78 3.6471 88.24 3.4118 82.35
Earth 4.0267 44.00 4.2319 52.17 4.1739 60.87
Biological Sciences 3.8800 48.00 3.6800 48.00 3.7867 32.00
Biomedical Sciences 4.0133 68.00 4.1333 72.00 3.7733 88.00
Health and Medicine 4.3733 56.00 3.7467 80.00 3.6800 52.00
Average 4.0873 64.18 4.0319 71.64 3.9149 72.72
A.2.4 Visual Supervised Fine-tuning
We fine-tuned the Qwen2-VL-2B model on our dataset for one epoch with LoRA (Hu et al., 2021), targeting all linear modules. We use the LLAMA-Factory framework for training (Zheng et al., 2024). The hyperparameters are provided in Table 9. The fine-tuning was conducted on a computing cluster with eight NVIDIA A100 GPUs, each with 40GB of memory, and the process took approximately 8 hours to complete.
Table 13: Hyperparameters for visual language pre-training on interleaved text and image data.
Hyperparameter Values
base model https://huggingface.co/meta-llama/Llama-2-7b-hfb
vision encoder https://huggingface.co/openai/clip-vit-large-patch14-336
projector 2-layer MLP
Stage 1: Projector Initialization
epochs $1$
global batch size 256
learning rate $0.001$
learning rate scheduler cosine
weight decay $0.0$
warmup ratio $0.03$
max length $4096$
tune LLM
tune vision encoder
tune projector
Stage 2: Visual Language Pre-training
epochs $1$
global batch size 128
learning rate $0.00005$
learning rate scheduler cosine
weight decay $0.0$
warmup ratio $0.03$
max length $4096$
tune LLM
tune vision encoder
tune projector
A.2.5 Visual Language Pre-training
In our case study experiments on the material generation task, we continuously pre-train a LLaMA2-7B model using our interleaved article and figure data to infuse more material science-relevant knowledge. Specifically, for pre-training on the interleaved text and image data, we follow the methodology outlined in (Lin et al., 2023).
Model Architecture
Following the approach outlined in (Liu et al., 2024; Lin et al., 2023), we extend the LLaMA2-7B model from a text-only model to a multimodal model by augmenting the LLM with a visual encoder to learn visual embeddings and a projector to bridge the embeddings between the text and visual modalities. Specifically, the visual encoder processes the image and outputs visual features. These features are then mapped into the word embedding space by the projector, creating visual tokens. These visual tokens are concatenated with the word tokens and fed into the LLM, allowing the model to integrate both text and visual information for generation. The specific LLM, visual encoder, and projectors used in our experiments are presented in Table 13.
Training Stages
The visual pre-training process (Lin et al., 2023) involves two stages:
  1. 1.
    Projection initialization: In this stage, the LLM and the visual encoder are both pre-trained and remain fixed. The projector, however, is randomly initialized. Only the projector is fine-tuned during this stage, using image-caption pairs from (Liu et al., 2024).
  2. 2.
    Visual language pre-training: During this stage, both the LLM and the projector are fine-tuned on the interleaved image and text data. This includes data from general domains provided by MMC4 (Zhu et al., 2024), as well as scientific articles and figures from our dataset MMSci. Previous research (Lin et al., 2023) has shown that tuning both the LLM and the projector yields better results than tuning only one of them. Throughout this stage, the visual encoder remains fixed.
We did not conduct the further visual instruction-tuning for this model, as our primary objective was to infuse scientific knowledge into the LLM for the consecutive text-only material generation task. The two stages were conducted on a computing cluster equipped with eight NVIDIA A100 GPUs, each with 40GB of memory. The first stage took approximately 4 hours, and the second stage took around 36 hours.
A.2.6 Materials Generation
As a case study to investigate whether scientific knowledge has been effectively infused into the LLM (LLaMA2-7B in our experiments) and whether it can enhance performance on material science-related tasks, we follow the methodology from (Gruver et al., 2024) to explore the material generation task. The primary objective is to format material crystal structures into text strings and fine-tuning the LLM to generate stable materials.
Prompt design
We adhere to the prompt design described in (Gruver et al., 2024). There are two types of prompts in the training data: the generation prompt with one or multiple conditions and infilling prompts, where partial crystal structure strings are masked and the model generates the masked parts. The specific prompt templates are shown below, adapted from (Gruver et al., 2024).
Generation Prompt Infilling Prompt
$<$s$>$Below is a description of a bulk material. [The chemical formula is Pm2ZnRh]. Generate a description of the lengths and angles of the lattice vectors and then the element type and coordinates for each atom within the lattice:
[ Crystal string ]$<$/s$>$
$<$s$>$Below is a partial description of a bulk material where one element has been replaced with the string “[MASK]”:
[ Crystal string with [MASK]s ]
Generate an element that could replace [MASK] in the bulk material:
[ Masked element ]$<$/s$>$
Blue text is the condition for generation. Purple text stands in for string encodings of atoms.
The formula condition as shown above is always included, while other conditions are sampled from the following: formation energy per atom, band gap, energy above hull, and space group number.
Evaluation
[…]
Metastability and stability are assessed based on the energy above the convex hull, denoted as $\hat{E}_{\text{hull}}$. Two approaches are employed to estimate $\hat{E}_{\text{hull}}$: M3GNet (Chen & Ong, 2022) and Density Functional Theory (DFT) using the VASP code (Hafner, 2008). For M3GNet, each sample undergoes relaxation using force and stress calculations before evaluating the energy of the final structure. For DFT, relaxation is performed using the VASP code, which provides more accurate results but requires significantly more computational resources. A material is considered metastable by M3GNet if the predicted energy above the hull, $E_{\text{hull}}^{\text{M3GNet}}$, is less than 0.1 eV/atom. Furthermore, if validated by DFT, the material must have $E_{\text{hull}}^{\text{DFT}}<0.0$ eV/atom to be considered stable. The percentages of such materials are reported over the total 10,000 inferences. We use the Materials Project (Jain et al., 2013) dated 2023-02-07.
Training Details
Following the approach in (Gruver et al., 2024), we utilize 4-bit quantization (Dettmers et al., 2021) and Low-Rank Adapters (LoRA) (Hu et al., 2021) for efficient fine-tuning. The model is trained with a batch size of 1 for 1 epoch. We set the LoRA rank to 8 and the LoRA alpha to 32. The learning rate is 0.0001, annealed by a cosine scheduler. The training was conducted on a single NVIDIA A100 GPU, took approximately 4 hours to complete.
Conditional Generation and Infilling Results
Due to space constraints, we did not include the results for the conditional materials generation and infilling tasks in the main paper. Here, we present these additional findings. The performance metrics reported are based on the same model used in the main paper. Our training data included two types of prompts: conditional generation prompts and infilling prompts. We compare our model LLaMA2-7B-MMSci, which has undergone continuous pre-training, with the original LLaMA2-7B that was trained without additional pre-training data. Both models were trained on datasets that included prompts for both conditional generation and infilling tasks under the same setup.
Following (Gruver et al., 2024), we performed 1,000 inferences for each condition in the conditional generation evaluation and 1,000 inferences for the infilling evaluation. For conditional generation evaluation, we assessed the percentage of generated materials that adhered to specified conditions, including formula, space group, and energy above the hull ($E_{\text{hull}}$). In the infilling evaluation, we measured diversity by computing the pairwise distance between generated samples and those from Matminer (Ward et al., 2018; Xie et al., 2021), focusing on composition and structure. Additionally, we evaluated metastability estimated by M3GNet. As seen in Table 14, LLaMA2-7B-MMSci, after continuous pre-training on our dataset MMSci, outperforms the original LLaMA2-7B across most metrics. This demonstrates its enhanced effectiveness in handling materials generation tasks.
Table 14: Evaluation of conditional materials generation and infilling tasks. Comp. Div. and Struct. Div. represent the composition and structure diversity, respectively. The two models are fine-tuned with the same training data and setup in our implementation.
Method Conditional Generation Infilling
[…]
LLaMA2-7B 0.85 0.14 0.58 10.60 0.16 64.20%
LLaMA2-7B-MMSci 0.87 0.22 0.59 8.31 0.52 77.74%
A.3 Datasheet
A.3.1 Motivation
With the advancement of large language and multimodal models, there is a growing demand for professional AI scientific assistants capable of comprehending and processing advanced, graduate-level scientific knowledge (noa, 2023; White, 2023; Vert, 2023). A crucial aspect of developing effective AI scientific assistants is their ability to understand academic scientific literature, which often includes complex figures such as data visualization plots, charts, schematic diagrams, macroscopic and microscopic photograph, and other specialized content from a variety of scientific fields. However, there is currently a lack of comprehensive evaluation for models’ understanding of advanced graduate-level multimodal scientific knowledge, especially in the context of complex figures across diverse scientific disciplines. Existing evaluations tend to focus on simpler charts and plots (Chen et al., 2020; Kahou et al., 2017; Siegel et al., 2016) and suffer from narrow scopes and lower quality (Li et al., 2024).
Our dataset, MMSci, is designed to address this gap. MMSci is a multimodal, multi-discipline dataset comprising high-quality, peer-reviewed articles and figures from 72 scientific disciplines, predominantly within the natural sciences. We created a benchmark to evaluate models’ understanding of graduate-level multimodal scientific knowledge across these disciplines. Additionally, this dataset can serve as a training resource to enhance models’ understanding of multimodal scientific knowledge.
A.3.2 Intended Use
This dataset is used to evaluate and enhance the large multimodal models (LVLMs)’ understanding of advanced multimodal scientific knowledge.
A.3.3 Data Collection
Data Source
The dataset comprises open-access articles published in Nature Communications [^7] https://www.nature.com/ncomms/. These articles are freely and permanently accessible upon publication under the Creative Commons Attribution 4.0 International (CC BY) License. Detailed information on the open-access policy of Nature Communications is available at https://www.nature.com/ncomms/open-access.
Data Collection Process
We collected various types of information for each article from the Nature Communications website. The articles’ information includes titles, abstracts, main body content, references, and PDF versions of the articles, all directly accessible from their respective sections on the article’s webpage (e.g., https://www.nature.com/articles/xxx, where “xxx” is the article’s unique ID). Additionally, figures and their captions were sourced from a dedicated figures section linked from each article’s main page (e.g., https://www.nature.com/articles/xxx/figures). This user-friendly platform facilitates easy acquisition of all necessary data, eliminating the needs for quality control and data filtering.
Annotations
The dataset does not include explicit annotations. Instead, the authors themselves carried out a small-scale manual review and classification of the image types specifically for analysis. No external annotators or crowdworkers were involved in this process.
Personal and Sensitive Information
The dataset does not include any personal or sensitive information. All article content is publicly accessible. All author information are also publicly available, and no personal information was explicitly extracted, stored, or used from the authors.
A.3.4 Social Impact and Ethical Considerations
Benefits
The benefits of our dataset are two-fold: (1) Evaluation Benchmark: This dataset serves as a valuable evaluation benchmark for assessing the understanding of large multimodal models (LVLMs) regarding scientific articles and figures. (2) Training Resources: It can be used as a training resource to enhance LVLMs’ understanding of scientific articles and figures, improving their performance in various scientific and research-related tasks.
Risks and Ethical Considerations
However, there are potential risks and ethical considerations to address: (1) Misuse in Academic Integrity: The advancement of AI research assistants facilitated by this dataset could potentially lead to misuse, such as academic fraud, fabrication, or improper assistance in academic work. We strongly encourage users to exercise caution and responsibility when using AI assistants, ensuring they are employed ethically and correctly. (2) Data Misinterpretation and Hallucination: There is a risk of misinterpreting the dataset’s content, leading to inaccurate conclusions or misuse of scientific information. Users should critically assess and validate the AI-generated outputs against established scientific knowledge and principles.
A.3.5 Limitations
Our dataset MMSci provides a comprehensive multimodal dataset across 72 scientific disciplines and serves as both a benchmark and a training resource. However, there are some limitations in our current exploration. (1) Due to limited resources, we were unable to evaluate a wide range of large-scale open-source LVLMs. (2) Our benchmark primarily assesses models’ understanding of scientific figures using the figures and captions. The dataset still provide other valuable resources that could be used to create additional tasks, such as single- and multimodal questions aimed at evaluating models’ scientific knowledge. We plan to explore these opportunities in future work. Despite these limitations, we believe MMSci will be a valuable resource for the research community. All data will be made publicly available.
A.3.6 Author Statement
The authors declare full responsibility for any rights violations, including but not limited to intellectual property rights and privacy rights, that may arise from the publication and use of this dataset. We confirm that all data provided is licensed under appropriate licenses, ensuring legal compliance and transparency.
A.3.7 Hosting, Licensing, and Maintenance Plan
The dataset will be hosted on GitHub, offering reliable and secure access. We commit to maintaining the repository with regular updates, security patches, and user support to ensure the data’s integrity and usability over time. Licensing terms will be clearly communicated to users, adhering to the appropriate data licenses to promote proper usage and distribution. The data is licensed under the CC BY 4.0 License, which permits sharing and adaptation with proper attribution. The primary codebase for our project is licensed under the Apache 2.0 License.
A.4 Examples
We present several figures as our case study to illustrate multiple-choice questions under three setting in Figure 7, 8, 9, respectively.
Figure 7: An example of the multi-choice questions (Fig2Cap). The example is within the material sciences subject, sourced from (Kang et al., 2014). The options include the correct main caption of the given figure and three main captions from other figures within the same article.
Figure 8: An example of the multi-choice questions (SubFig2Cap). The example is within the biochemistry subject, sourced from (Wang et al., 2016).
Figure 9: An example of the multi-choice questions (SubCap2Fig). The example is within the cancer subject, sourced from (da Silva-Coelho et al., 2017).